jennifer is 3 years older than Luke and 2 years younger than Evan.Their total age is 44. How old is each child?
Answer by jorel555(1290)

Let J equal jennifer's age, L=luke's age, E=evan's age:
J+L+E=44
J-3=Luke
E-2=Jennifer
J+(J-3)+(J+2)=44
3J-1=44
3J=45
J=15
Jennifer is 15, Luke is 12, and Evan is 17!!!!!!